OpenStudy (anonymous):

A bag contains 50 marbles, 10 of which are blue, 8 are red, 20 are green, and 12 are purple. Marcie takes a marble out of the bag, records the color, and returns it to the bag. How many red marbles should she expect after 500 trials? 80 100 120 200

OpenStudy (anonymous):

A ?

OpenStudy (anonymous):

well.i would do it cuz i subtracted 50 minus everything ...so yea i would go with tht :)

OpenStudy (anonymous):

OKay thank you

OpenStudy (welshfella):

probability of drawing a red marble on one draw = 8/50 = 4/25

OpenStudy (welshfella):

expected number of reds for 500 trials = 4/25 * 500 = 4*20 = 80